OpenStudy (ivancsc1996):

I'm taking an advanced single variable calculus class and I need help with this problem: Is there a number n in the naturals that satisfies the following inequality? (The inequality is written below)

OpenStudy (ivancsc1996):

\[n \in \mathbb{N} \rightarrow \sum_{1}^{n}\frac{ 1 }{ 1+i ^{2} }\ge \frac{ \pi }{ 2 } ? \]

OpenStudy (ivancsc1996):

I am supposed to not know anything about convergence and approximation of sums through integrals *wink wink*

ganeshie8 (ganeshie8):